A 1.1-kg object is suspended from a vertical spring whose spring constant is 170 N/m. (a) Find the amount by which the spring is stretched from its unstrained length. (b) The object is then pulled straight down by an additional distance of 0.10 m and released from rest. Find the speed with which the object passes through its original position on the way up.
Answer #1

Somo At equilibrium, the spring is equal to the weight Fs = Fg Krang x = mg - (lukg) (9.20 m/s2) K 170N/m- p = 0.063) (6) By